Circular No. 7244

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999dn IN NGC 7714 K. Ayani, Bisei Astronomical Observatory; R. Furusho, Institute of

Physical and Chemical Research; H. Kawakita, Gunma Astronomical Observatory; M. Fujii, Fujii Bisei Observatory; and H. Yamaoka, Kyushu University, report: “A low-dispersion spectrum (400–750 nm) of SN 1999dn (cf. IAUC 7241) was obtained with the Bisei 1.01-m telescope on Aug. 21.7 UT. The spectrum resembles that of SN 1997X posted at http://oir-www.

harvard.edu/cfa/oir/Research/supernova/spectra/sn97x.gif, which

indicates that this supernova is of type Ic shortly after maximum light (though it is also similar to the spectra of type-Ia supernovae two weeks past maximum, lacking however a deep, broad absorption feature at 500 nm). Our spectrum reveals reddish continuum with strong absorption features whose minima are at 620 nm (Si II), 564 nm (Na I D and He I blend), and 468 nm (He I). These wavelengths are converted to the rest frame of the host galaxy, using observed v_{r} = 2790 km/s, which is determined by the Hα and Hβ emission lines. Narrow Na I D interstellar lines are weak or absent. The expansion velocity derived from the Si II absorption minimum is 7300 km/s.”

Y.-l. Qiu, Beijing Astronomical Observatory (BAO), writes: “A lowdispersion spectrum (range 460–800 nm) obtained by L. Huang, B. Yao, and H. Li on Aug. 21.8 UT with the BAO 2-m telescope (+ BFOSC spectrograph) suggests that SN 1999dn is a type-Ia supernova. A deep, broad blueshifted absorption of Si II is seen with a minimum at 618.5 nm, which implies an expansion velocity of 8020 km/s. A narrow Hα emission line from the host galaxy NGC 7714 gives a redshift of 2650 km/s. The Na D line and Fe II at 553.5 nm are strong. The overall spectrum is similar to that of SN 1989B, two weeks after maximum. Since this supernova was not seen 11 days before discovery (cf. IAUC 7241), it is supposed that the supernova is just past optical maximum, and its spectrum has perhaps evolved faster than those of normal type-Ia supernovae.”

M. Turatto, L. Rizzi, M. Salvo, and E. Cappellaro, Padua; S. Benetti, Telescopio Nazionale Galileo; and F. Patat, European Southern Observatory, note that a spectrum (350–900 nm; resolution 0.1 nm) of SN 1999dn, obtained with the 1.5-m Danish telescope (+ DFOSC) at La Silla on Aug. 25.3 UT, is similar to those of the type-Ic SNe 1994I and 1996aq near maximum light, with broad lines of Ca II (H, K, and infrared triplet), Si II, and Na I with P-Cyg profiles. The continuum is similar to that of SN 1996aq but bluer than that of SN 1994I. There is no evidence of interstellar absorption. The expansion velocity (from Ca II lines) is 13 000 km/s.

1999 August 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green